OpenStudy (abbygirl15):

Multiply. Express your answer in lowest terms. 1/6 times 5/7 I don't get? please help!

OpenStudy (anonymous):

5/42 is the answer

OpenStudy (anonymous):

@BerzerkC Please don't just give direct answers.

OpenStudy (anonymous):

well, how do u give indirect answers for these kinds of problems?

OpenStudy (anonymous):

When you multiply fractions, you can multiply straight across (numerator x numerator and denominator x denominator). So your problem would go like this... \[\frac{ 1 }{ 6 }\times \frac{ 5 }{ 7 }=\frac{ 1\times5 }{ 6\times7 }\] Then simplify the top and simplify the bottom.

OpenStudy (anonymous):

@BerzerkC like that

OpenStudy (anonymous):

well fine multipy the 2 numerators with the 2 denominators to get the answer

OpenStudy (anonymous):

ok thanks!

OpenStudy (abbygirl15):

ok both of you get medals.

OpenStudy (anonymous):

thanks!